# Uncertainty-Aware PPG-2-ECG for Enhanced Cardiovascular Diagnosis using Diffusion Models

### v1

@@ E.2 Training Details for the Conversion Model
Concerning training, the model architecture is trained on MIMIC-III. We randomly split the data into 70% for training and 30% for validation after preprocessing. The model architecture is trained for 10,000 epochs, employing a batch size of 2,048 PPG and ECG recordings across 8$\times$×V100 GPUs. The noise schedule progresses linearly over 1,000 time steps, starting at 1e-6 and ending at 1e-2. For optimization, we employed the standard Adam optimizer with a fixed learning rate of 1e-4. An $L_{2}$ loss is utilized to estimate the noise in the input instance.
For evaluation, we sample $K=100$ ECG candidates for each PPG signal using the diffusion-based framework. We adopt a non-stochastic diffusion process as proposed in [28], with skip jumps of 10 iterations, resulting in a total of $T=1000/10=100$ diffusion iterations.
E.3 Training Details for the Classification Model
@@
As discussed earlier, the presence of each cardiovascular condition in the CinC dataset is imbalanced, and thus CinC poses significant challenges for developing a robust classifier. In Figure 6 (red) we report the frequency of each cardiovascular condition in the dataset. To address this issue, we implement a novel class-balance sampling technique. Given the multi-label nature of the classification, where class dependencies exist, our strategy not only ensures a uniform class distribution for positive samples but also carefully selects negative samples from a balanced pool reflecting common class co-occurrences. This approach significantly aids in achieving a more uniform label distribution, the effectiveness of which is demonstrated in Figure 6 (blue).
The classification architecture utilized in our studies is based on a simplified VGG [46] model with 8 layers, incorporating batch normalization modules to promote stability, along with dropout at ratio of 0.2 to address overfitting. Similar to the conversion model, we modify the VGG architecture for 1D input signals by reconfiguring the convolutional layers for 1D operation.
Optimization is performed on 1$\times$×V100 GPU using the Adam optimizer and an $L_{2}$ weight decay of 1e-4. The model is trained over 20 epochs with a batch size of 128, employing a learning rate of 1e-3 with a linear decay and using binary cross-entropy loss as the loss function.
@@
Appendix F Classification Strategies
In our experiments we assess several classification strategies, all relying on paired ECG and PPG signals and their corresponding labels, as obtained using MIMIC-III [38] and CinC [35]. We assume the availability of pretrained classifiers, $f_{\mathcal{X}}:\mathcal{X}\rightarrow\mathbb{R}^{+}$ and $f_{\mathcal{Y}}:\mathcal{Y}\rightarrow\mathbb{R}^{+}$, for input ECG and PPG signals, respectively. These classifiers output classification likelihoods considering cardiovascular conditions.
@@ Appendix F Classification Strategies
Appendix G Exploring the Validity of Synthesized PPG Signals within MIMIC-III
Figure 7: MIMIC-III Results: (a) ROC Curves illustrating classification performance. (b) Risk-Coverage curves [25] demonstrating uncertainty quantification performance. (c) Incremental classification error as a function of the number of ECG solutions utilized for each PPG signal.
The main question arising from the results depicted in Section 5.2 is whether using synthesized PPG signals for drawing the various conclusions is a fair and trust-worthy strategy. One potential medical rationale for this approach is that PPG signals inherently encompass noise. Consequently, employing synthesized PPG signals, even if noisy themselves, can accurately model real-world scenarios. Still, in order to address this question in a more direct form, we return to the MIMIC-III [38] dataset, this time using a much smaller held-out subset of it that contains binary labels indicating the presence of atrial fibrillation (AFib). Thus, for this heart condition, we can evaluate classification performance for both true PPG signals and synthesized ones.
As previously stated in Section 5.2, the MIMIC-III matched waveform database lacks cardiovascular labels. However, textual cardiac reports are available for a subset of patients within MIMIC-III, referred to as the MIMIC-III clinical database [38]. In order to extract AFib labels from this subset, we filtered 400 patients, with 200 reported cases of AFib and 200 with normal sinus rhythm. The determination of AFib or sinus rhythm presence was made by searching for the terms "atrial fibrillation" and "sinus rhythm" within the cardiac reports. We note that extracting labels using this searching approach may provide erroneous labels.
@@ Appendix G Exploring the Validity of Synthesized PPG Signals within MIMIC-III
Focusing on the main theme of this experiment, we see that the performance of classifying the original PPG (ORANGE) is comparable to the alternative of operating on the synthetic signals (RED). In terms of AUROC, the performances are nearly identical. When evaluating the area under the Risk-Coverage curve [25], we observe a similar trend; the uncertainty performance of both approaches is almost identical when considering around half of the test samples. These findings offer empirical evidence supporting the validity of our reversal conversion model.
Appendix H CinC Numerical Results
@@
Table 3: CinC Results: Mean and standard error values of macro-level AUROC and AURC across three random seeds, shown for each cardiovascular condition examined. See Appendix F for the classification strategies considered.
@@
1) Ventricular premature beat Original ECG 0.8953 $\pm$ 0.0081 0.0042 $\pm$ 0.0005
Synthesized PPG 0.6724 $\pm$ 0.0033 0.0216 $\pm$ 0.0009
Synthesized ECG: SSC Mean 0.6290 $\pm$ 0.0194 0.0171 $\pm$ 0.0006
Synthesized ECG: SSC Random 0.6344 $\pm$ 0.0087 0.0178 $\pm$ 0.0014
Synthesized ECG: ESC (Ours) 0.8006 $\pm$ 0.0072 0.0077 $\pm$ 0.0006
2) Atrial flutter Original ECG 0.9710 $\pm$ 0.0004 0.0056 $\pm$ 0.0001
Synthesized PPG 0.8457 $\pm$ 0.0018 0.0289 $\pm$ 0.0006
Synthesized ECG: SSC Mean 0.7648 $\pm$ 0.0170 0.0424 $\pm$ 0.0040
Synthesized ECG: SSC Random 0.8018 $\pm$ 0.0035 0.0380 $\pm$ 0.0009
Synthesized ECG: ESC (Ours) 0.8816 $\pm$ 0.0009 0.0220 $\pm$ 0.0003
3) Sinus rhythm Original ECG 0.9728 $\pm$ 0.0004 0.0175 $\pm$ 0.0003
Synthesized PPG 0.8111 $\pm$ 0.0010 0.1309 $\pm$ 0.0005
Synthesized ECG: SSC Mean 0.7638 $\pm$ 0.0018 0.1700 $\pm$ 0.0012
Synthesized ECG: SSC Random 0.7773 $\pm$ 0.0009 0.1618 $\pm$ 0.0010
Synthesized ECG: ESC (Ours) 0.8547 $\pm$ 0.0004 0.1114 $\pm$ 0.0013
4) Atrial fibrillation Original ECG 0.9489 $\pm$ 0.0008 0.0059 $\pm$ 0.0001
Synthesized PPG 0.8263 $\pm$ 0.0007 0.0283 $\pm$ 0.0005
Synthesized ECG: SSC Mean 0.8320 $\pm$ 0.0144 0.0262 $\pm$ 0.0004
Synthesized ECG: SSC Random 0.8582 $\pm$ 0.0044 0.0317 $\pm$ 0.0010
Synthesized ECG: ESC (Ours) 0.9196 $\pm$ 0.0016 0.0147 $\pm$ 0.0003
5) Supraventricular premature beats Original ECG 0.9080 $\pm$ 0.0017 0.0068 $\pm$ 0.0002
Synthesized PPG 0.6673 $\pm$ 0.0033 0.0302 $\pm$ 0.0011
Synthesized ECG: SSC Mean 0.6176 $\pm$ 0.0223 0.0313 $\pm$ 0.0027
Synthesized ECG: SSC Random 0.6404 $\pm$ 0.0054 0.0351 $\pm$ 0.0006
Synthesized ECG: ESC (Ours) 0.7739 $\pm$ 0.0069 0.0163 $\pm$ 0.0006
6) Bradycardia Original ECG 0.8072 $\pm$ 0.0224 0.0010 $\pm$ 0.0002
Synthesized PPG 0.7734 $\pm$ 0.0154 0.0023 $\pm$ 0.0001
Synthesized ECG: SSC Mean 0.6405 $\pm$ 0.0284 0.0143 $\pm$ 0.0057
Synthesized ECG: SSC Random 0.6927 $\pm$ 0.0098 0.0028 $\pm$ 0.0005
Synthesized ECG: ESC (Ours) 0.7853 $\pm$ 0.0113 0.0011 $\pm$ 0.0001
7) Pacing rhythm Original ECG 0.9734 $\pm$ 0.0048 0.0009 $\pm$ 0.0002
Synthesized PPG 0.7258 $\pm$ 0.0119 0.0054 $\pm$ 0.0001
Synthesized ECG: SSC Mean 0.6148 $\pm$ 0.0212 0.0139 $\pm$ 0.0010
Synthesized ECG: SSC Random 0.6193 $\pm$ 0.0181 0.0137 $\pm$ 0.0008
Synthesized ECG: ESC (Ours) 0.7966 $\pm$ 0.0061 0.0055 $\pm$ 0.0002
8) Sinus tachycardia Original ECG 0.9932 $\pm$ 0.0005 0.0015 $\pm$ 0.0001
Synthesized PPG 0.9674 $\pm$ 0.0003 0.0094 $\pm$ 0.0002
Synthesized ECG: SSC Mean 0.9451 $\pm$ 0.0030 0.0136 $\pm$ 0.0010
Synthesized ECG: SSC Random 0.9263 $\pm$ 0.0030 0.0184 $\pm$ 0.0009
Synthesized ECG: ESC (Ours) 0.9792 $\pm$ 0.0018 0.0050 $\pm$ 0.0005
9) 1st degree AV block Original ECG 0.9634 $\pm$ 0.0011 0.0029 $\pm$ 0.0001
Synthesized PPG 0.6350 $\pm$ 0.0043 0.0365 $\pm$ 0.0006
Synthesized ECG: SSC Mean 0.6139 $\pm$ 0.0092 0.0332 $\pm$ 0.0009
Synthesized ECG: SSC Random 0.6184 $\pm$ 0.0050 0.0347 $\pm$ 0.0011
Synthesized ECG: ESC (Ours) 0.7193 $\pm$ 0.0095 0.0203 $\pm$ 0.0010
10) Sinus bradycardia Original ECG 0.9935 $\pm$ 0.0002 0.0025 $\pm$ 0.0001
Synthesized PPG 0.9438 $\pm$ 0.0008 0.0304 $\pm$ 0.0005
Synthesized ECG: SSC Mean 0.8932 $\pm$ 0.0028 0.0565 $\pm$ 0.0028
Synthesized ECG: SSC Random 0.9005 $\pm$ 0.0010 0.0502 $\pm$ 0.0011
Synthesized ECG: ESC (Ours) 0.9580 $\pm$ 0.0003 0.0199 $\pm$ 0.0011
11) Sinus arrhythmia Original ECG 0.9512 $\pm$ 0.0003 0.0039 $\pm$ 0.0001
Synthesized PPG 0.7741 $\pm$ 0.0032 0.0207 $\pm$ 0.0004
Synthesized ECG: SSC Mean 0.6999 $\pm$ 0.0039 0.0272 $\pm$ 0.0019
Synthesized ECG: SSC Random 0.7344 $\pm$ 0.0037 0.0268 $\pm$ 0.0006
Synthesized ECG: ESC (Ours) 0.8771 $\pm$ 0.0006 0.0086 $\pm$ 0.0001
@@
Table 4: CinC Results: Mean and standard error values of AUROC and AURC [25] across three random seeds, shown for each cardiovascular condition examined. See Appendix F for the classification strategies considered.
This section extends the empirical classification results referring to CinC [35], that were outlined in Section 5.2. We provide here the numerical results of the mean metric values along with their corresponding standard errors.
Tables 3 and 4 display the AUROC (Area Under the ROC curve) and AURC (Area Under the Risk-Coverage curve [25]) metrics. Specifically, Table 4 outlines AUROC/AURC mean and standard error values associated with each classification strategy per cardiovascular condition, whereas Table 3 provides macro-level metrics derived by averaging over all examined cardiovascular conditions.

### v3

@@ D.2 Training Details for the Conversion Model
Concerning training, the model architecture is trained on MIMIC-III. We split the data into training and validation sets after preprocessing. Specifically, we randomly select 400 patients for the validation set while the remaining patients use for training. For each patient in the training set we split the raw PPG/ECG signals into pairs of 10-seconds (length 1024) signals, resulting in 15,225,389 PPG/ECG signals. For each patient in the validation set, we sample 100 distinct 10-seconds (length 1024) PPG/ECG signals from the raw signals, resulting in 40,000 PPG/ECG signals.
The model architecture is trained for 10,000 epochs, employing a batch size of 2,048 PPG and ECG recordings across 8$\times$×V100 GPUs. The noise schedule progresses linearly over 1,000 time steps, starting at 1e-6 and ending at 1e-2. For optimization, we employed the standard Adam optimizer with a fixed learning rate of 1e-4. An $L_{2}$ loss is utilized to estimate the noise in the input instance.
For evaluation, we sample $K=100$ ECG candidates for each PPG signal using the diffusion-based framework, and repeat the generation process for 3 different seeds. We adopt a non-stochastic diffusion process as proposed in Song et al. (2020), with skip jumps of 10 iterations, resulting in a total of $T=1000/10=100$ diffusion iterations.
D.3 Training Details for the Classification Model
@@
As discussed earlier, the presence of each cardiovascular condition in the CinC dataset is imbalanced, and thus CinC poses significant challenges for developing a robust classifier. In Figure 9 (red) we report the frequency of each cardiovascular condition in the dataset. Note that the challenge in sampling class-balanced batches arises from the fact that samples may be positive for multiple cardiovascular conditions at once. To address this issue, we implement a novel class-balance sampling technique. Given the multi-label nature of the classification, where class dependencies exist, our strategy not only ensures a uniform class distribution for positive samples but also carefully selects negative samples from a balanced pool reflecting common class co-occurrences. This approach significantly aids in achieving a more uniform label distribution, the effectiveness of which is demonstrated in Figure 9 (blue).
Specifically, we consider a training set $S_{\text{Train}}$, where each pair $(X,C)\in S_{\text{Train}}$ satisfies $X\in\mathcal{X}$ and $C\in\{0,1\}^{L_{\text{max}}}$. Here, $C$ is a binary vector where $C_{i}=1$ if and only if the label $i$ is present in $X$. At each timestep, we first sample a random label uniformly from the $L_{\text{max}}=11$ cardiovascular conditions. We then ensure that both positive $(X^{p},C^{p})$ and negative $(X^{n},C^{n})$ samples for this condition are included in the batch. Additionally, we control the ratio of the presence of the major label, denoted by $1\leq l\leq L_{\text{max}}$, as one condition is often predominant, representing the ”normal” condition. In our case, this label is sinus rhythm (see Figure 9). We set this ratio to be $p_{l}=0.2$. For a formal summary of this method, please refer to Algorithm 3.
Algorithm 3 Class-Balanced Batch Sampling
1:Training set $S_{\text{ Train}}\subseteq\mathcal{X}\times\mathcal{C}$ s.t. $\mathcal{C}=\{0,1\}^{L_{\text{max}}}$. Batch size $2b\in\mathbb{N}$. Major label index $1\leq l\leq L_{\text{max}}$. Major label ratio $p_{l}\in[0,1]$.
2:$S_{\text{ Batch}}\leftarrow\{\}$
3:for $i=1$ to $b$ do
4:     $L\sim U(\{1,2,\dots,L_{\text{max}}\})$
5:     $(X^{p},C^{p})\sim U(\{(X,C)\in S_{\text{ Train}}:C_{L}=1\})$
6:     if $L=l$ then $(X^{n},C^{n})\sim U(\{(X,C)\in S_{\text{ Train}}:C_{L}=0\})$
7:     else $(X^{n},C^{n})\sim U(\{(X,C)\in S_{\text{ Train}}:C_{L}=0,C_{l}\sim B(p_{\text{l}})\})$
8:     $S_{\text{ Batch}}\leftarrow\{(X^{p},C^{p}),(X^{n},C^{n})\}$
9:end for
10:Class-balanced batch $S_{\text{ Batch}}$.
The classification architecture utilized in our studies is based on a simplified VGG (Simonyan & Zisserman, 2014) model with 8 layers, incorporating batch normalization modules to promote stability, along with dropout at ratio of 0.2 to address overfitting. Similar to the conversion model, we modify the VGG architecture for 1D input signals by reconfiguring the convolutional layers for 1D operation.
The training process is implemented above CinC using a cross-validation scheme with 3 seeds, where for each seed, we create a separated 80-20 training and validation set patient split after preprocessing. Specifically, our training sets contain 65,535 10-seconds (length 1024) ECG signals, while the validation sets contain 16,386 10-seconds (length 1024) ECG signals.
Optimization is performed on 1$\times$×V100 GPU using the Adam optimizer and an $L_{2}$ weight decay of 1e-4. The model is trained over 20 epochs with a batch size of 128, employing a learning rate of 1e-3 with a linear decay and using binary cross-entropy loss as the loss function. Since the primary focus of our paper is on the conversion task rather than the classification tasks, which are used for validation and to motivate the consideration of uncertainty in the conversion, the decision thresholds were simply fixed to 0.5 as standard.
Appendix E Classification Strategies
In our experiments we assess several classification strategies, all relying on paired ECG and PPG signals and their corresponding labels, as obtained using MIMIC-III (Johnson et al., 2016) and CinC (Reyna et al., 2021). We assume the availability of pretrained classifiers, $f_{\mathcal{X}}:\mathcal{X}\rightarrow\mathbb{R}^{+}$ and $f_{\mathcal{Y}}:\mathcal{Y}\rightarrow\mathbb{R}^{+}$, for input ECG and PPG signals, respectively. These classifiers output classification likelihoods considering cardiovascular conditions.
@@ Appendix E Classification Strategies
Note that neither RDDM (Shome et al., 2023) nor CardioGAN (Sarkar & Etemad, 2021) provides publicly available code, which are currently state-of-the-art conversion models. This limits our ability to generate ECG signals using their methods and evaluate their classification performance. However, our evaluation setup is more comprehensive, making CardioGAN and RDDM comparable to the evaluated classification strategies. Both approaches combine the MMSE estimate with a random single sample, since they both rely on a single random ECG solution that tends toward the mean due to mode collapse, as shown in Table 1.
Appendix F Exploring the Validity of Synthesized PPG Signals
The main question arising from the results depicted in Section 5.2 is whether using synthesized PPG signals for drawing the various conclusions is a fair and trust-worthy strategy. One potential medical rationale for this approach is that PPG signals inherently encompass noise. Consequently, employing synthesized PPG signals, even if noisy themselves, can accurately model real-world scenarios. Still, in order to address this question in a more direct form, we return to the MIMIC-III (Johnson et al., 2016) dataset, which contains true PPG signals, and conduct three additional studies described hereafter.
F.1 Achieving Similar Trends with True PPG Signals and True Labels
Figure 10: MIMIC-III Results: (a) ROC Curves illustrating classification performance. (b) Risk-Coverage curves (Geifman & El-Yaniv, 2017) demonstrating uncertainty quantification performance. (c) Incremental classification error as a function of the number of ECG solutions utilized for each PPG signal.
First, we use a much smaller held-out subset of the MIMIC-III database that contains binary labels indicating the presence of atrial fibrillation (AFib). Therefore, for this heart condition, we can evaluate classification performance for both true PPG signals and synthesized ones.
As previously stated in Section 5.2, the MIMIC-III matched waveform database lacks cardiovascular labels. However, textual cardiac reports are available for a subset of patients within MIMIC-III, referred to as the MIMIC-III clinical database (Johnson et al., 2016). In order to extract AFib labels from this subset, we filtered 400 patients, with 200 reported cases of AFib and 200 with normal sinus rhythm. The determination of AFib or sinus rhythm presence was made by searching for the terms ”atrial fibrillation” and ”sinus rhythm” within the cardiac reports. We note that extracting labels using this searching approach may provide erroneous labels.
@@ F.1 Achieving Similar Trends with True PPG Signals and True Labels
Focusing on the main theme of this experiment, we see that the performance of classifying the original PPG (ORANGE) is comparable to the alternative of operating on the synthetic signals (RED). In terms of AUROC, the performances are nearly identical. When evaluating the area under the Risk-Coverage curve (Geifman & El-Yaniv, 2017), we observe a similar trend; the uncertainty performance of both approaches is almost identical when considering around half of the test samples. These findings offer empirical evidence supporting the validity of our reversal conversion model.
F.2 Assessing the High Quality of Synthetic PPG Signals
Syn PPG 1-FD ↓ 100-FD ↓
DDIM (T=25) 3.8365 $\pm$ 0.0021 3.8248 $\pm$ 0.0013
DDIM (T=50) 0.4049 $\pm$ 0.0013 0.3927 $\pm$ 0.0007
DDIM (T=100) 0.1181 $\pm$ 0.0011 0.1036 $\pm$ 0.0002
Table 3: MIMIC-III Results: Quality assessment of our reversed diffusion-based conversion model. The results show the mean metric estimates and their standard errors across the three seeds.
Second, we conduct an experiment that evaluates the signal quality of the reversed conversion model over MIMIC-III. Specifically, given the ECG-2-PPG algorithm, we generate many PPG signals; we then compute the distance between the distributions of the true PPG and the synthesized ones. The results are presented in Table 3, which is similar to Table 1, showing (with T=100) that the distributions are nearly indistinguishable from each other (with almost zero FD). Furthermore, the quality of the synthesized PPG signals improves with more diffusion iterations.
F.3 Assessing the High Quality of Synthetic ECG Signals from Synthetic PPG Ones
UA-P2E 1-FD ↓ 100-FD ↓
Syn ECGs from True PPG 0.3198 $\pm$ 0.0020 0.2379 $\pm$ 0.0005
Syn ECGs from Syn PPG 0.3940 $\pm$ 0.0023 0.3164 $\pm$ 0.0005
Table 4: MIMIC-III Results: Quality assessment of our synthetic ECG signals derived from synthetic PPG vs those derived from true PPG. The results show the mean metric estimates and their standard errors across the three seeds.
Finally, we conduct an experiment that testifies the signal quality after cycle transformation: ECG to PPG to ECG. Specifically, we sample (synthetic) PPG signals from true ECG signals in MIMIC-III, then re-sample (synthetic) ECG signals from these synthetic PPG ones. Finally, we measure their distribution distance compared to the true ECG signals. The results are presented in Table 4, which is also similar to Table 1. These results indicate that the synthetic ECG signals derived from synthetic PPG are almost as high-quality as those derived from true PPG. The negligible loss in quality is expected, given that the synthetic PPG signals are not perfect. This empirical evidence further supports the use of synthetic PPG data in our work.
Appendix G Supplementary Experimental Analysis of CinC Results
This section extends the empirical classification results referring to CinC (Reyna et al., 2021), that were outlined in Section 5.2.
G.1 Reported Numerical Results
@@
Table 5: CinC Results: Mean and standard error values of macro-level AUROC and AURC across three random seeds. See Appendix E for the classification strategies considered.
Here, we provide the numerical results of the mean metric values along with their corresponding standard errors reported in Figure 3 and Figure 4.
Tables 5 and 7 display the AUROC (Area Under the ROC curve) and AURC (Area Under the Risk-Coverage curve (Geifman & El-Yaniv, 2017)) metrics. Specifically, Table 7 outlines AUROC/AURC mean and standard error values associated with each classification strategy per cardiovascular condition, whereas Table 5 provides macro-level metrics derived by averaging over all examined cardiovascular conditions.
G.2 Evaluating Additional Metrics
Classification Strategy Macro-TPR ↑ Macro-TNR ↑ Macro-F1 ↑
Original ECG 0.6684 $\pm$ 0.0009 0.9802 $\pm$ 0.0002 0.6679 $\pm$ 0.0005
Synthesized PPG 0.2850 $\pm$ 0.0007 0.9694 $\pm$ 0.0000 0.3139 $\pm$ 0.0003
Synthesized ECG: SSC Mean 0.2542 $\pm$ 0.0081 0.9560 $\pm$ 0.0012 0.2663 $\pm$ 0.0081
Synthesized ECG: SSC Random 0.3218 $\pm$ 0.0008 0.9509 $\pm$ 0.0000 0.3100 $\pm$ 0.0009
Synthesized ECG: ESC (Ours) 0.2902 $\pm$ 0.0039 0.9783 $\pm$ 0.0002 0.3154 $\pm$ 0.0017
Table 6: CinC Results: Mean and standard error values of macro-level sensitivity (TPR), specificity (TNR), and F1 scores, across three random seeds. See Appendix E for the classification strategies considered.
To enhance clinical relevance and facilitate comparisons with future studies, we evaluate additional metrics related to the classification performance in mitigating class imbalance. Specifically, we evaluate sensitivity (TPR), specificity (TNR), and F1 scores across different classification strategies. In Table 6 we provide the reported results.
It is evident from Table 6 that our classification approach outperforms in specificity (TNR) compared to sensitivity (TPR). The random ECG signal classification approach, stems from the generation process of our diffusion model, demonstrates better sensitivity than multi-solution approach. We attribute this phenomenon to the rarity of various cardiovascular conditions in the CinC dataset (see Figure 8), which makes it challenging to achieve high sensitivity, even with the original ECG signals where no PPG signals are considered.
While our approach achieves superior performance in specificity, it can be used in clinical setting as a better detector for the absence of cardiovascular conditions on average, while a single ECG sample of our diffusion model is evident to be better on the detection of the presence of the cardiovascular conditions on average.
@@ G.2 Evaluating Additional Metrics
1) Ventricular premature beat Origin ECG 0.8953 $\pm$ 0.0081 0.0042 $\pm$ 0.0005
Syn PPG 0.6724 $\pm$ 0.0033 0.0216 $\pm$ 0.0009
Syn ECG: SSC Mean 0.6290 $\pm$ 0.0194 0.0171 $\pm$ 0.0006
Syn ECG: SSC Random 0.6344 $\pm$ 0.0087 0.0178 $\pm$ 0.0014
Syn ECG: ESC (Ours) 0.8006 $\pm$ 0.0072 0.0077 $\pm$ 0.0006
2) Atrial flutter Origin ECG 0.9710 $\pm$ 0.0004 0.0056 $\pm$ 0.0001
Syn PPG 0.8457 $\pm$ 0.0018 0.0289 $\pm$ 0.0006
Syn ECG: SSC Mean 0.7648 $\pm$ 0.0170 0.0424 $\pm$ 0.0040
Syn ECG: SSC Random 0.8018 $\pm$ 0.0035 0.0380 $\pm$ 0.0009
Syn ECG: ESC (Ours) 0.8816 $\pm$ 0.0009 0.0220 $\pm$ 0.0003
3) Sinus rhythm Origin ECG 0.9728 $\pm$ 0.0004 0.0175 $\pm$ 0.0003
Syn PPG 0.8111 $\pm$ 0.0010 0.1309 $\pm$ 0.0005
Syn ECG: SSC Mean 0.7638 $\pm$ 0.0018 0.1700 $\pm$ 0.0012
Syn ECG: SSC Random 0.7773 $\pm$ 0.0009 0.1618 $\pm$ 0.0010
Syn ECG: ESC (Ours) 0.8547 $\pm$ 0.0004 0.1114 $\pm$ 0.0013
4) Atrial fibrillation Origin ECG 0.9489 $\pm$ 0.0008 0.0059 $\pm$ 0.0001
Syn PPG 0.8263 $\pm$ 0.0007 0.0283 $\pm$ 0.0005
Syn ECG: SSC Mean 0.8320 $\pm$ 0.0144 0.0262 $\pm$ 0.0004
Syn ECG: SSC Random 0.8582 $\pm$ 0.0044 0.0317 $\pm$ 0.0010
Syn ECG: ESC (Ours) 0.9196 $\pm$ 0.0016 0.0147 $\pm$ 0.0003
5) Supraventricular premature beats Origin ECG 0.9080 $\pm$ 0.0017 0.0068 $\pm$ 0.0002
Syn PPG 0.6673 $\pm$ 0.0033 0.0302 $\pm$ 0.0011
Syn ECG: SSC Mean 0.6176 $\pm$ 0.0223 0.0313 $\pm$ 0.0027
Syn ECG: SSC Random 0.6404 $\pm$ 0.0054 0.0351 $\pm$ 0.0006
Syn ECG: ESC (Ours) 0.7739 $\pm$ 0.0069 0.0163 $\pm$ 0.0006
6) Bradycardia Origin ECG 0.8072 $\pm$ 0.0224 0.0010 $\pm$ 0.0002
Syn PPG 0.7734 $\pm$ 0.0154 0.0023 $\pm$ 0.0001
Syn ECG: SSC Mean 0.6405 $\pm$ 0.0284 0.0143 $\pm$ 0.0057
Syn ECG: SSC Random 0.6927 $\pm$ 0.0098 0.0028 $\pm$ 0.0005
Syn ECG: ESC (Ours) 0.7853 $\pm$ 0.0113 0.0011 $\pm$ 0.0001
7) Pacing rhythm Origin ECG 0.9734 $\pm$ 0.0048 0.0009 $\pm$ 0.0002
Syn PPG 0.7258 $\pm$ 0.0119 0.0054 $\pm$ 0.0001
Syn ECG: SSC Mean 0.6148 $\pm$ 0.0212 0.0139 $\pm$ 0.0010
Syn ECG: SSC Random 0.6193 $\pm$ 0.0181 0.0137 $\pm$ 0.0008
Syn ECG: ESC (Ours) 0.7966 $\pm$ 0.0061 0.0055 $\pm$ 0.0002
8) Sinus tachycardia Origin ECG 0.9932 $\pm$ 0.0005 0.0015 $\pm$ 0.0001
Syn PPG 0.9674 $\pm$ 0.0003 0.0094 $\pm$ 0.0002
Syn ECG: SSC Mean 0.9451 $\pm$ 0.0030 0.0136 $\pm$ 0.0010
Syn ECG: SSC Random 0.9263 $\pm$ 0.0030 0.0184 $\pm$ 0.0009
Syn ECG: ESC (Ours) 0.9792 $\pm$ 0.0018 0.0050 $\pm$ 0.0005
9) 1st degree AV block Origin ECG 0.9634 $\pm$ 0.0011 0.0029 $\pm$ 0.0001
Syn PPG 0.6350 $\pm$ 0.0043 0.0365 $\pm$ 0.0006
Syn ECG: SSC Mean 0.6139 $\pm$ 0.0092 0.0332 $\pm$ 0.0009
Syn ECG: SSC Random 0.6184 $\pm$ 0.0050 0.0347 $\pm$ 0.0011
Syn ECG: ESC (Ours) 0.7193 $\pm$ 0.0095 0.0203 $\pm$ 0.0010
10) Sinus bradycardia Origin ECG 0.9935 $\pm$ 0.0002 0.0025 $\pm$ 0.0001
Syn PPG 0.9438 $\pm$ 0.0008 0.0304 $\pm$ 0.0005
Syn ECG: SSC Mean 0.8932 $\pm$ 0.0028 0.0565 $\pm$ 0.0028
Syn ECG: SSC Random 0.9005 $\pm$ 0.0010 0.0502 $\pm$ 0.0011
Syn ECG: ESC (Ours) 0.9580 $\pm$ 0.0003 0.0199 $\pm$ 0.0011
11) Sinus arrhythmia Origin ECG 0.9512 $\pm$ 0.0003 0.0039 $\pm$ 0.0001
Syn PPG 0.7741 $\pm$ 0.0032 0.0207 $\pm$ 0.0004
Syn ECG: SSC Mean 0.6999 $\pm$ 0.0039 0.0272 $\pm$ 0.0019
Syn ECG: SSC Random 0.7344 $\pm$ 0.0037 0.0268 $\pm$ 0.0006
Syn ECG: ESC (Ours) 0.8771 $\pm$ 0.0006 0.0086 $\pm$ 0.0001
@@
Table 7: CinC Results: Mean and standard error values of AUROC and AURC (Geifman & El-Yaniv, 2017) across three random seeds, shown for each cardiovascular condition examined. See Appendix E for the classification strategies considered.
Appendix H Why the Original ECG Performance Sets an Upper Bound for PPG-based Approach?
The performance of the original ECG signal represents an upper bound for classification performance, including our own method. Intuitively, this discrepancy arises because cardiovascular labels are directly linked to heart conditions determined by ECG signals, while PPG signals only capture blood volume fluctuations, which provide partial information about the heart. As a result, PPG signals are a degraded version of ECG signals, leading to a natural decline in performance due to the loss of information. Even when generating ECG candidates from PPG signals, some information loss is unavoidable, which further reduces performance. Our approach mitigates this issue by accounting for conversion uncertainty, yielding superior results compared to other PPG-based strategies.
From a more theoretical perspective, we adopt a Markovian framework: $Y\rightarrow X\rightarrow C$, where $Y$ represents the PPG signal, $X$ denotes the ECG signal, and $C$ is the cardiovascular condition label. Under this assumption, $\pi(C|X,Y)=\pi(C|X)$, meaning that, given the ECG signal $X$, the PPG signal $Y$ does not provide any additional information about the cardiovascular condition $C$. Consequently, when we degrade the class information from $X$ to $Y$, performance naturally declines due to the inherent loss of information.